OpenStudy (anonymous):

Square root of 5x-6=x

OpenStudy (akshay_budhkar):

\[\sqrt{5x-6}=x\]

OpenStudy (akshay_budhkar):

square on both the sides

OpenStudy (akshay_budhkar):

you get x^2=5x-6

OpenStudy (akshay_budhkar):

therefore x^2-5x+6=0

OpenStudy (akshay_budhkar):

therefore x^2-3x-2x+6=0

OpenStudy (akshay_budhkar):

therefore x(x-3)-2(x-3)=0

OpenStudy (anonymous):

My choices are 2/3 , 2 , 3 , and no solution .

OpenStudy (akshay_budhkar):

therefore x=3 or x=2

OpenStudy (akshay_budhkar):

can it be multiple?

OpenStudy (akshay_budhkar):

if yes then 2 and 3

OpenStudy (anonymous):

As long as they are positive, np